P2-xp-q=0 solve by charpit method


1
Expert's answer
2022-01-03T16:18:42-0500

Given differential equation is "p^2 - xp-q=0\Rightarrow p^2 = q + xp" (1)

We can write it as "f(x,y,z,p,q) = q + xp - p^2 = 0"


"\frac{\partial f }{\partial x } = p" , "\frac{\partial f }{\partial y } = 0" , "\frac{\partial f }{\partial z } = 0" , "\frac{\partial f }{\partial p} = x-2p" , "\frac{\partial f }{\partial q } = 1"


Hence, equation will be


"\frac{dx}{-\frac{\partial f }{\partial p } } = \frac{dy}{-\frac{\partial f }{\partial q } } = \frac{dz}{-p\frac{\partial f }{\partial p }-q\frac{\partial f }{\partial q } } = \frac{dp}{\frac{\partial f }{\partial x }+p\frac{\partial f }{\partial z } } = \frac{dz}{\frac{\partial f }{\partial y }+q\frac{\partial f }{\partial z } }"


"\frac{dx}{2p-x } = \frac{dy}{-1 } = \frac{dz}{-p(x-2p)-q } = \frac{dp}{p} = \frac{dz}{0 }"


Taking "\frac{dy}{-1 } = \frac{dp}{p}"

Integrating it we get,

"p = ae^{-y}"

Then from given equation (1),

"q = -axe^{-y} + a^2e^{-2y}"


Now, "dz = pdx + qdy = ae^{-y}dx + (-axe^{-y}+a^2e^{-2y})dy"


Integrating it both sides, we get


"z = axe^{-y}-\frac{1}{2}a^2e^{-2y} + b"